When Sheila asked Margie to deliever camp supplies across the lake, they agreed to meet on the water. Shelia's boat can travel at 10km/h, while Margie's boat can do 12km/h. If the lake is 6.6 km wide and Sheila and Margie leave at the same time, how much time will it be before they meet?
Answer by funmath(2933)

Distance = rate*time or d=rt
They travel the same amount of time: t (in hours)
Sheila travels: 10t (km)
Margie travels: 12t (km)
Altogether they travel: 6.6 (km)
Our problem to solve is:
10t+12t=6.6
(10+12)t=6.6
22t=6.6
22t/22=6.6/22
t=.3 (hours)
Most of us don't know how long that really is, so convert it to minutes:
t=.3(60)min
t=18 min